OpenStudy (anonymous):

Find the 9th term of the arithmetic sequence. a7 = -42 and a11 = -28; a9 = ____.

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

@awks

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let d be the common difference

jimthompson5910 (jim_thompson5910):

a8 = a7 + d

jimthompson5910 (jim_thompson5910):

a9 = a8 + d a9 = a7 + d + d a9 = a7 + 2d

jimthompson5910 (jim_thompson5910):

a10 = a9 + d a10 = a7 + 2d + d a10 = a7 + 3d see the pattern?

OpenStudy (anonymous):

hmm sorta walk me thru it

jimthompson5910 (jim_thompson5910):

what would a11 be?

OpenStudy (anonymous):

mm 14

jimthompson5910 (jim_thompson5910):

how are you getting 14

OpenStudy (anonymous):

idk i guessed i have no clue

jimthompson5910 (jim_thompson5910):

do you see how I got a8 = a7 + d ?

OpenStudy (anonymous):

hmm u added 1

jimthompson5910 (jim_thompson5910):

I took a7 and I added d

jimthompson5910 (jim_thompson5910):

if you know a7 and d, you can find a8

jimthompson5910 (jim_thompson5910):

then notice I wrote a9 = a8 + d

OpenStudy (anonymous):

ahh :/

jimthompson5910 (jim_thompson5910):

see how I'm getting that?

OpenStudy (anonymous):

continue

jimthompson5910 (jim_thompson5910):

so if you have any term, you add d to it to get the next term

jimthompson5910 (jim_thompson5910):

say I have a10 how do I get a11 ?

OpenStudy (anonymous):

-35

jimthompson5910 (jim_thompson5910):

if you have any term, you add d to it to get the next term

jimthompson5910 (jim_thompson5910):

what do you do to a10 to get a11?

jimthompson5910 (jim_thompson5910):

any ideas?